The mean score for all indviduals who took the Law School Admission Test (LSAT) 10 2010 was 155 and the standard deviation was 7.
What is the z score for a person, who received an LSAT socre of 167"?
a. +1.71
b. +1.64
c. +1.74
d. -1.71
What is the raw score for a person with a z score of -0.86?
a. 164
b. 146
c. 149
d. 161

Answer by jim_thompson5910(35256)


Part 1) What is the z score for a person, who received an LSAT score of 167?


z = (x-mu)/sigma


z = (167-155)/7


z = 12/7


z = 1.71428571428571 this is approximate


z = 1.71 rounding to two decimal places


So the answer is choice A) +1.71


--------------------------------------------------------------------------------


Part 2) What is the raw score for a person with a z score of -0.86?


z = (x-mu)/sigma


-0.86 = (x-155)/7


-0.86*7 = x-155


-6.02 = x-155


-6.02+155 = x


148.98 = x


x = 148.98


x = 149 rounding to the nearest whole number


So the answer is choice C) 149